500^3=s^3
The volume of a cube is given by V=s^3 where the length of a side , Find the length of a side of a cube if the volume is 500cm^3. Round answer to three decimal places

Answer by edjones(8007)

s^3=v
s^3=500
s=500^(1/3)
=7.937 cm
.
Ed